OpenStudy (anonymous):

Choose the correct slope and y-intercept of the line indicated by this linear equation. 4y - 8x + 12 = 0

OpenStudy (anonymous):

y=mx+c 4y-8x+12=0 4y=8x-12 y=(8x-12)/4 y=2x-3 from the equation y=mx+c m=slope=2 and the intercept,c=-3